Ramos is concerned about the number of prescriptions his elderly clients have. He would like to create a 99% confidence interval for the mean number of prescriptions per client with a maximum error of 2 prescriptions. Assuming a standard deviation of 5.2 prescriptions, what is the minimum number of clients he must sample?
1
Expert's answer
2020-09-21T15:44:33-0400

solution


maximum error "=2"

Standard deviation"=5.2"


Confidence level "=99\%"

Hence "\alpha = 1\%"

The Z-score "Z_{\alpha/2}=Z_{0.005}=2.58"


Sample size (n):



"n = (\frac{Z_{\alpha/2}* Standard\ Deviation}{Maximum\ Error})^2"

"= (\frac{2.58*5.2}{2})^2=44.997"

answer: the minimum number of clients required for the test is 45